what are the asymptotes of 2tanx?
Answer by Alan3354(69443)

what are the asymptotes of 2tanx?
-----------
x = pi/2 + n*pi, n = 0,1,2,3...
x = pi/2 + n*pi, n = 0,-1,-2,-3...
My +/- symbol no longer works.